Circular No. 5620

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1737 II = 1862 III = 1992t)

A message from H. Kosai, National Astronomical Observatory, Tokyo, reports the discovery of a comet by Tsuruhiko Kiuchi, with the suggestion that it might be P/Swift-Tuttle with perihelion time in mid-December. Confirmation of the discovery by several observers in Canada, the U.S. and Japan leaves no doubt that this identification with the Perseid parent comet is correct. The identification in turn confirms the suggestion (Lynn 1902, Obs. 25, 304; Marsden 1973, A.J. 78, 662; see also IAUC 5330 and 5586) that Kegler’s 1737 observations were indeed of P/Swift-Tuttle.

1992 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.26.75694 11^{h}47.^{m}5  +59^{o}00′  11.5 Kiuchi
27.22465 11 50 34.^{s}71 +59 04 35′′.3  Tatum
27.41424 11 51 45.90 +59 05 26.6 12.5 Sugie
27.42326 11 51 49.22 +59 05 33.1  _{”}
27.43368 11 51 53.49 +59 05 34.1  _{”}
27.44792 11 51 58.70 +59 05 37.0  Tatum
27.49897 11 52.23 +59 05.5 Scotti

T. Kiuchi (Usuda, Nagano). 25 × 150 binoculars. Comet diffuse without

condensation, diameter 4′. J. B. Tatum (University of Victoria). 0.25-m Schmidt. A. Sugie (Dynic Astronomical Observatory). 0.25-m f/3.4 Schmidt. Strong

condensation. Communicated by S. Nakano. J. V. Scotti (Kitt Peak). 0.91-m Spacewatch telescope. Well-condensed

nucleus, fan-shaped coma > 3′.4 wide, tail to southwest.

Precise linkage of the observations, even at two apparitions, is not possible without the involvement of large nongravitational forces. The following orbital elements, by the undersigned, are a compromise that fit the current observations exactly but leave large discordances in 1862:

Epoch = 1992 Dec. 4.0 TT

T = 1992 Dec. 12.391 TT  ω = 152.^{o}979
e = 0.96362  Ω = 139.430 }2000.0
q = 0.95876 AU  i = 113.408
a = 26.35441 AU n^{o} = 0.007285 P = 135.29 years

1992 September 27 Brian G. Marsden